OpenStudy (marcelie):

help please number 22

OpenStudy (marcelie):

evaluate the integral

OpenStudy (marcelie):

@agent0smith

OpenStudy (marcelie):

@zepdrix

OpenStudy (shawn):

seems like a trig sub

OpenStudy (marcelie):

yes , but having trouble

OpenStudy (marcelie):

so would it be x= tan x dx: sec^2 deta

OpenStudy (shawn):

x = tan(t) would be more appropriate

OpenStudy (marcelie):

oh hmm would it tan deta ?

OpenStudy (shawn):

it's just a dummy variable. theta is fine if you'd like to be consistent with the book

OpenStudy (marcelie):

oh okay so then would it be like this ? \[\int\limits_{}^{}\sqrt{(\tan \theta)^2 +1}\]

OpenStudy (usukidoll):

isn't it sec^2x-1 = tan^2x?

OpenStudy (marcelie):

oh yes sorry

OpenStudy (shawn):

what about dx?

OpenStudy (marcelie):

\[\int\limits_{}^{}\sqrt{(\sec^2\theta-1)+1} d \theta \]

zepdrix (zepdrix):

That was sort of a silly way to apply your Pythagorean Identity :) Yes, \(\rm \sec^2\theta-1=\tan^2\theta\) but that means \(\rm \sec^2\theta=\tan^2\theta+1\)

zepdrix (zepdrix):

Oh oh, woops you messed something up with your differential I think.

OpenStudy (marcelie):

D:

zepdrix (zepdrix):

\[\large\rm x=\tan \theta\]\[\large\rm dx=\sec^2\theta~d \theta\]

zepdrix (zepdrix):

\[\large\rm \int\limits \sqrt{\color{royalblue}{x}^2+1}~\color{orangered}{dx}\quad=\int\limits \sqrt{\color{royalblue}{\tan}^2\color{royalblue}{\theta}+1}~\color{orangered}{\sec^2\theta~d \theta}\]

zepdrix (zepdrix):

And then you apply your Pythagorean Identity,\[\large\rm =\int\limits \sqrt{\sec^2\theta}~\sec^2\theta~d \theta\]

OpenStudy (marcelie):

\[\int\limits_{}^{}(\sqrt{1+\tan^2\theta}) ( 1+\tan^2\theta)\]

zepdrix (zepdrix):

What? Why turn them back into tangents? That causes a problem for us. Now we have the +1 under the root again. The whole point of using a trig sub... is to get rid of the addition under the root. It makes it much easier to take a square root from that point.

OpenStudy (usukidoll):

oy... no switchy switchy back to tanx

OpenStudy (marcelie):

aah lol thats where i get stuck D:

zepdrix (zepdrix):

I meant that I was apply Pythagorean Identity from here,\[\large\rm \int\limits \sqrt{\tan^2\theta+1}~\sec^2\theta~d \theta\]to here,\[\large\rm \int\limits \sqrt{\sec^2\theta}~\sec^2\theta~d \theta\]

zepdrix (zepdrix):

Now.......... take square root.

OpenStudy (marcelie):

hmm wait question what happen to the +1 ?

zepdrix (zepdrix):

Our trigonometric Pythagorean Identity gets rid of it. \(\large\rm tan^2\theta+1=sec^2\theta\) There is no +1 on the right side of this equation, ya? We're replacing the tan^2 + 1 with the right side of the equation in our integral.

OpenStudy (marcelie):

oh wait nvm i see it .. it was from the identity right ?

OpenStudy (marcelie):

oh okay i hope i did it right \[\int\limits_{}^{} \sec^2\theta* \sec ^2 \theta \] d deta

zepdrix (zepdrix):

You didn't apply square root.

zepdrix (zepdrix):

\(\large\rm \sqrt{x^2}\ne x^2\)

OpenStudy (marcelie):